# Probabilistic estimates of the diameters of the Rubik’s Cube groups

### v4

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from any initial configurations. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probabilistic theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. It is shown that the probabilistically estimated diameter is approximated by $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ I Objective
The diameter of the Cayley graph of the Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) group (hereafter the diameter) is the fewest number of turns to solve the Cube from any initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
@@ II Background
Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf,Korf (1983, 1985) especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theory.Zassenhaus (1982); Fiat et al. (1989); Joyner (2008); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) Furthermore, its similarities with problems appearing in quantum mechanics,Corli et al. (2021) statistical mechanics,Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014) high-energy physics,Czech et al. (2011) and cryptographyDiaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problemGumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probabilistic theory (see Appendix A).
@@ IV The 2$\times$2$\times$2 Cube
One of the eight corner cubies (say, the front-top-left one) of the 2$\times$2$\times$2 Cube is held fixed as the orientation anchor of the whole Cube (here, a “cubie” stands for a 1$\times$1$\times$1 constituent cube). The remaining seven cubies’ positions can be permuted in $7!$ ways and their orientations can take $3^{7}$ ways, except the last (seventh) cubie’s orientation is fixed by the first six (it is saidHofstadter (1981) that the “sanity” of the seven cubies is maintained by only one of three orbits). Therefore, the total number of configurations $N$ is
@@ IV.2 Quarter-turn metric
Hence, there is a negligible chance that the predicted diameter is actually 13 (not 14), but a small (14%) probability that it is 15. In this case, the predicted most likely (86%) value of the diameter (14) proves to be exact. See Sec. VIII.3 for our analysis of the metric-dependence of the accuracy of the probabilistic estimation.
@@ IV.3 Semi-quarter-turn metric
In this newly introduced metric, only three (3) clockwise 90 turns ($R$, $D$, and $B$) are allowed in each step. With an explicit enumeration, we find
@@ IV.4 Bi-quarter-turn metric
In this another new metric, all nine (9) turns in the half-turn metric plus six (6) additional bi-quarter turns ($RD$, $R^{-1}D^{-1}$, $DB$, $D^{-1}B^{-1}$, $BR$, $B^{-1}R^{-1}$) are elementary operations. This metric is intentionally highly redundant.
@@ V The 3$\times$3$\times$3 Cube
The first fraction denotes the number of ways in which the eight corner cubies can be positioned ($8!$) and oriented ($3^{8}$) divided by the three orbits, only one of which maintains the sanity.Hofstadter (1981) The second fraction is the number of ways in which the twelve edge cubies can be placed ($12!$) and oriented ($2^{12}$) divided by the four orbits, only one of which has the sanity.Hofstadter (1981) The expected number of random configurations to generate all $N$ configurations is therefore
@@ VIII Discussion
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Actual Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (101) (this work).
@@ VIII.1 Comparison between predicted and actual results
For the 3$\times$3$\times$3 Cube, Rokicki et al.Rokicki et al. (2013) gave the actual and estimated numbers (“positions”) of new configurations generated at each step. They are compared with our predicted values in Fig. 5. In this case, the rising part and the peak of the actual curve is accurately reproduced by the prediction, but the falling part decays too slow in the latter, overestimating the diameter by two, although the Rokicki et al.’s values in the last five steps are approximate. The actual curve has a rapid falloff of $e^{-22.3t}$, although it likely decreases faster than exponentially. On the other hand, the predicted curve decays exponentially as $e^{-13.3t}$ corresponding to $r=13.33$. It again suggests that our probabilistic logic tends to underestimate the accerelated rate of exhausting remaining unrealized configurations in the final few steps. However, the exact agreement of the predicted and actual diameters in the quarter-turn metric may contradict this view, given the smaller value of $r=9.37$ and even slower decay of the predicted curve in this metric.
@@ VIII.2 Size dependence
With the branching ratio $r$ and the number of configurations $N$, both of which can be relatively easily determined as has been done in the foregoing sections, the probabilistic estimate of the diameter (which must be distinguished from the correct diameter) is approximated by
$\frac{\ln N}{\ln r}+\frac{\ln N}{r}.$ (101)
@@
VIII.3 Metric dependence
Figure 6: Percentage contributions of the types of turns in the shortest solutions of the 2$\times$2$\times$2 Cube for the half-turn metric. Column “Total” plots the cumulative contributions of turns in all of the solutions. Turns are executed in the same order as the histogram is stacked, i.e., $R$, $R^{2}$, $R^{-1}$, $D$, $D^{2}$, $D^{-1}$, $B$, $B^{2}$, $B^{-1}$.
Figure 7: Same as Fig. 6 for the quarter-turn metric.
Figure 8: Same as Fig. 6 for the semi-quarter-turn metric. “T” stands for “Total.”
Figure 9: Same as Fig. 6 for the bi-quarter-turn metric.
Not much attention seems to have been paid to the metric dependence of the diameter. If one assumes that all nine turns ($X$, $X^{2}$, and $X^{-1}$, where $X=R$, $D$, $B$) of the 2$\times$2$\times$2 Cube in the half-turn metric are equally useful to solve all configurations, the diameter in the quarter-turn metric (the correct value is 14) should be calculable from the same (11) in the half-turn metric as
$\frac{11}{9}\times 6+\frac{11}{9}\times 3\times 2=14.7,$ (102)
where the second term doubles the number of turns for each of the three $X^{2}$ compound turns. It slightly overestimates the correct diameter of 14. The diameter (19) in the semi-quarter-turn metric is predicted as
$\frac{11}{9}\times 3+\frac{11}{9}\times 3\times 2+\frac{11}{9}\times 3\times 3=22.0,$ (103)
which is too large by three. The diameter (26) of the 3$\times$3$\times$3 Cube in the quarter-turn metric is obtainable from the same (20) in the half-turn metric as
$\frac{20}{18}\times 12+\frac{20}{18}\times 6\times 2=26.7,$ (104)
which is again overestimated slightly. These overestimations suggest that $X^{2}$ (180) turns are less used especially near the diameter.
This is borne out in Fig. 6, where the numbers of types of turns in the half-turn metric used in the shortest solutions are shown as histograms. While the contributions of nine turns are comparable with one another in all the solutions (in the rightmost column “Total”), the contribution of $X^{2}$ is less than those of $X$ and $X^{-1}$ in the last few steps of the diameter. In particular, the contribution of $R$ becomes progressively more important near the diameter. This is primarily because $R$ happens to be the first turn applied in our computer program, and other subsequent turns have greater probabilities of generating redundant configurations as the remaining unrealized configurations become fewer.
One might expect that the more uneven the contributions of different types of turns are, the less accurate our probabilistic estimates of the diameters become, since the latter are oblivious of the different types of turns. This expectation is not necessarily supported by our data in Figs. 69. The most erroneous prediction for the 2$\times$2$\times$2 Cube is in the semi-quarter-turn metric, but this metric has one of the most uniform usage of its three turns (Fig. 8). In contrast, Fig. 9 shows that the bi-quarter-turn metric is superfluous with the $90^{\circ}$ turns less and less used near the diameter. However, the error in the predicted diameter is less than that in the semi-quarter-turn metric. The most accurate prediction is for the quarter-turn metric, but the use of different turns seems no more uniform (Fig. 7) than in the other metrics. It may be concluded that the present probabilistic estimation is largely insensitive to the metrics.

### v5

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from the hardest initial configuration. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probabilistic theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. The probabilistically estimated diameter is shown to obey the approximate semi-analytical formula of $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ I Objective
The diameter of the Cayley graph of the Rubik’s CubeHofstadter (1981); Singmaster (1981); Biggs (1993); Joyner (2008); Ferenc (2024); West (2019) group (hereafter the diameter) is the fewest number of turns to solve the Cube in the hardest initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
@@ II Background
Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf,Korf (1983, 1985) especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theoryZassenhaus (1982); Fiat et al. (1989); Joyner (2008); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) and graph theory.Beenker and Vanlint (1988); Biggs (1993); Ferrero and Hanusch (2014); West (2019) Furthermore, its similarities with problems appearing in quantum mechanics,Corli et al. (2021) statistical mechanics,Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014) high-energy physics,Czech et al. (2011) and cryptographyDiaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problemGumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probabilistic theory (see Appendix A). See pages 26 and 34 of Singmaster’s book,Singmaster (1981) where the author develops a similar argument. It is also related to mathematical models of epidemicsHill (2020) and network theory.Newman (2008)
@@ IV The 2$\times$2$\times$2 Cube
One of the eight corner cubies (say, the front-top-left one) of the 2$\times$2$\times$2 Cube is held fixed as the orientation anchor of the whole Cube (here, a “cubie” stands for a 1$\times$1$\times$1 constituent cube). The remaining seven cubies’ positions can be permuted in $7!$ ways and their orientations can take $3^{7}$ ways, except the last (seventh) cubie’s orientation is fixed by the first six (it is saidHofstadter (1981) that the “sanity” of the seven cubies is maintained by only one of three orbitsJoyner (2008)). Therefore, the total number of configurations $N$ (or the order of the 2$\times$2$\times$2 Rubik’s Cube group) is
@@ IV.2 Quarter-turn metric
Hence, there is a negligible chance that the predicted diameter is actually 13 (not 14), but a small (14%) probability that it is 15. In this case, the predicted most likely (86%) value of the diameter (14) proves to be exact.
@@ IV.3 Semi-quarter-turn metric
In this newly introduced metric, only three (3) clockwise 90 turns ($R$, $D$, and $B$) are allowed in each step. Since the inverses are not included, its “group” may not meet the mathematical definition of a group and the corresponding graph may not be a Cayley graph. Nevertheless, all of $N$ configurations can be reached by combinations of these three turns and the “diameter” is still well defined.
With an explicit enumeration, we find
@@ IV.4 Bi-quarter-turn metric
In this another new metric, all nine (9) turns in the half-turn metric plus six (6) additional bi-quarter turns ($RD$, $R^{-1}D^{-1}$, $DB$, $D^{-1}B^{-1}$, $BR$, $B^{-1}R^{-1}$) are elementary operations. This metric is intentionally highly redundant. Lacking some inverses, its “group” and “Cayley graph” may not meet the proper mathematical definitions, but the “diameter” is still well defined.
@@ V The 3$\times$3$\times$3 Cube
The first fraction denotes the number of ways in which the eight corner cubies can be positioned ($8!$) and oriented ($3^{8}$) divided by the three orbits,Joyner (2008) only one of which maintains the sanity.Hofstadter (1981) The second fraction is the number of ways in which the twelve edge cubies can be placed ($12!$) and oriented ($2^{12}$) divided by the four orbits,Joyner (2008) only one of which has the sanity.Hofstadter (1981) The expected number of random configurations to generate all $N$ configurations is therefore
@@ VIII Discussion
Cube Metric $r$ [^1] Branching ratio. See Eq. (4). Actual Predicted [^2] Probabilistic estimate based on a modified version of the coupon collector’s problem (this work). Predicted [^3] Equation (101) unless otherwise noted (this work).
@@ VIII.1 Comparison between predicted and actual results
For the 3$\times$3$\times$3 Cube, Rokicki et al.Rokicki et al. (2013) gave the actual and estimated numbers (“positions”) of new configurations generated at each step. They are compared with our predicted values in Fig. 5. In this case, the rising part and the peak of the actual curve is accurately reproduced by the prediction, but the falling part decays too slow in the latter, overestimating the diameter by two, although the Rokicki et al.’s values in the last five steps are approximate. The actual curve has a rapid falloff of $e^{-22.3t}$, although it likely decreases faster than exponentially. On the other hand, the predicted curve decays exponentially as $e^{-13.3t}$ corresponding to $r=13.33$. It again suggests that our probabilistic logic tends to underestimate the accelerated rate of exhausting remaining unrealized configurations in the final few steps. However, the exact agreement of the predicted and actual diameters in the quarter-turn metric may contradict this view, given the smaller value of $r=9.37$ and even slower decay of the predicted curve in this metric.
Singmaster,Singmaster (1981) using the same argument, inferred that nearly all configurations should be generated at 17.3 steps, which is only slightly less than our value or the one of Rokicki et al.Rokicki et al. (2013) On this basis, Singmaster conjectured that “every position can be achieved in at most 20 moves,” which is in exact agreement with the correct value determined by Rokicki et al.Rokicki et al. (2013) some 33 years later, although no account was given about the addition of 3 moves.
@@ VIII.2 Size dependence
With the branching ratio $r$ and the number of configurations $N$, both of which can be relatively easily determined as has been done in the foregoing sections, the probabilistic estimate ($d$) of the diameter (which must be distinguished from the correct diameter) is approximated by
$d\approx\frac{\ln N}{\ln r}+\frac{\ln N}{r}.$ (101)